# Not All Adapters Matter: Selective Adapter Freezing for Memory-Efficient Fine-Tuning of Language Models

### v1

@@ Abstract
Transformer-based large-scale pre-trained models achieve great success, and fine-tuning, which tunes a pre-trained model on a task-specific dataset, is the standard practice to utilize these models for downstream tasks. Recent work has developed adapter-tuning, but these approaches either still require a relatively high resource usage. Through our investigation, we show that each adapter in adapter-tuning does not have the same impact on task performance and resource usage. Based on our findings, we propose SAFE, which gradually freezes less-important adapters that do not contribute to adaptation during the early training steps. In our experiments, SAFE reduces memory usage, computation amount, and training time by 42.85%, 34.59%, and 11.82%, respectively, while achieving comparable or better performance compared to the baseline. We also demonstrate that SAFE induces regularization effect, thereby smoothing the loss landscape.
@@ 1 Introduction
Large-scale pre-trained language models (PLMs) have manifested superior performance in various tasks Kenton and Toutanova (2019); Liu et al. (2019); Radford et al. (2019); Yang et al. (2019). However, training a PLMs from the scratch is usually time-consuming and resource-intensive. Common practice has been hence to fine-tune the large-scale pre-trained models by adapting all the parameters with the downstream tasks, i.e., full parameter fine-tuning (full-tuning).
Recently, Parameter-Efficient Fine-Tuning (PEFT), which focuses on optimizing a small fraction of parameters for downstream tasks, is receiving much attention Houlsby et al. (2019); Lester et al. (2021); Li and Liang (2021); Liu et al. (2022, 2023). Among various PEFT strategies, adapter-tuning has emerged as a prevalent method. It integrates lightweight modules, termed adapters, into each layer of PLMs and only tunes the adapters with the downstream tasks. As shown in Figure 1(a), the adapter-tuning methods Houlsby et al. (2019); Pfeiffer et al. (2021); Zaken et al. (2022); Hu et al. (2021); Zhang et al. (2022), significantly reduce the number of trainable parameters, compared to the full-tuning, while exhibiting better performance for a downstream task.
@@ 2 Related Work
Figure 2: We quantify (a) accuracy and (b) memory usage of adapter-tuning by injecting an adapter into each transformer layer of $\text{BERT}_{\text{base}}$ model on MNLI and QNLI dataset from GLUE.
@@
Resource Efficient Fine-Tuning: Several works have tried to target resource efficient fine-tuning. AdapterDrop Rücklé et al. (2021), randomly excludes partial adapters from each training step. However, it cannot de-allocate the activation memory for the adapters, because of the random selections — an adapter excluded from training in a step can be included in training in the following steps. SparseAdapterHe et al. (2022) applies unstructured pruning to the adapters. However, it also does not reduce the actual memory usage — this is because the weight matrices pruned with zero values still need to be fully allocated in the memory. LoRAPruneZhang et al. (2023) employs structured pruning for LoRA. Unfortunately, the aforementioned methods usually have an adverse impact on the accuracy. MEFT Liao et al. (2024) applies a reversible model to PEFT. By using the reversable network, MEFT calculates activations with accumulated outputs of layers, without saving the intermediate activations reducing the activation memory. However, calculations of the activations severely degrades the training time performance.
@@ 3 Motivation
Figure 3: We visualize the representation similarity of the trained model and the model during training in adapter-tuning of the $\text{BERT}_{\text{base}}$ model on datasets from GLUE.
@@ 4.2 Adapter Freezing
In the freezing stage, SAFE gradually freezes adapters based on their importance score. At $t_{w}$-th epoch, SAFE compares the importance score of adapters with threshold $\tau_{T}$. If the importance score of an adapter is lower than $\tau_{T}$, SAFE identifies the adapter as a freezing candidate. After identifying freezing candidates, SAFE gradually freezes them based on a moving threshold $\tau_{T}$ until $t_{f}$-th epoch — it increases 0 to $\tau_{T}$ [^2] We empirically determine $\tau_{T}$ and final freezing epochs $t_{f}$ based on extensive experiments with various models and datasets. between $t_{w}$-th and $t_{f}$-th epochs following a cubic schedule as Zhang et al. (2022):
@@
where $t$ is the current epoch, $t_{w}$ is the number of initial warm-up epochs, $t_{f}$ is the number of final freezing epochs. This ensures that a large number of adapters participate in adaptation during the early stages of training, allowing SAFE to more accurately identify important adapters. As training progresses, less important adapters are gradually frozen, with different training periods assigned to each adapter based on their importance.
@@ 4.3 Regularization Effect of SAFE
By selectively freezing less critical adapters, SAFE induces a regularization effect within the model. In transformer-based PLM $\mathcal{N}_{0}$, each of the $l$ transformer blocks $T_{l}$ is equipped with a distinct set of parameters $\theta^{0}_{l}$ for $l\in\{1,\ldots,n\}$. To reduce the computational overhead of directly fine-tuning all parameters $\theta^{0}_{l}$, lightweight adapters $\Delta\theta_{l}$ are introduced. To clarify how introducing adapters contributes to performance enhancements, Fu et al. Fu et al. (2023) formalize the optimization function as follows:
@@
where $\theta=\theta^{0}+M\Delta\theta$ and $M\in\{0,1\}^{m\times m}$, with $m=\text{dim}(\theta)$, serves as a diagonal matrix for selective parameter adjustment. Each diagonal element $M_{ii}\in\{0,1\}$ indicates whether the corresponding parameter of $\Delta\theta_{i}$ is active ($1$) or inactive (00), with all off-diagonal elements $M_{ij}$ set to 0. The regularization term is crucial for explaining how parameter constraints introduced by adapters can enhance model performance on downstream tasks. The $rank(M)$ is bounded by $m$, reflecting full capacity for parameter adaptation within each transformer block. However, such an approach can lead to excessive computation. In contrast, our study explores the implications of constraining the $rank(M)$ to a reduced upper limit of $\frac{m}{l}$ by selectively activating $\Delta W_{l}$ for $T_{l}$. This constraint not only optimizes computational efficiency but also preserves the adaptability essential for superior performance on downstream tasks, as evidenced by our empirical results detailed in Section 5.3.
@@ 5.1 Experimental Setting
Baselines: To evaluate the effectiveness of SAFE, we benchmark against state-of-the-art PEFT method, LoRA Hu et al. (2021). We compare SAFE with two effective sparse training methods, AdapterDrop Rücklé et al. (2021) and SparseAdapter He et al. (2022). SAFE’s performance is further compared with other PEFT methods such as Houlsby Houlsby et al. (2019), Pfeiffer Pfeiffer et al. (2020), BitFit Zaken et al. (2022), and the adaptive method AdaLoRA Zhang et al. (2022) to demonstrate its versatility and applicability across different adapter-tuning frameworks (Figure 1, Appendix B). Comprehensive details on the experimental setup and hyperparameters, such as training epochs and batch sizes, can be found in Appendix C.
@@ 5.2.1 Natural Language Understanding
Compared to AdapterDrop, SAFE provides up to 2.69% higher score (MRPC) while reducing memory usage by 49.47% (0.91% higher GLUE score and 40.47% reduced memory usage on average). This is because AdapterDrop reduces computation costs (FLOPs) by randomly dropping adapters for each step, whereas SAFE selectively freezes less-important adapters preserving critical adapters and thus achieving better performance. AdapterDrop also does not lower memory usage because the memory allocated to the dropped adapters cannot be de-allocated for the next step — adapters dropped in a step may not be dropped in the next step. Compared to SparseAdapter, SAFE reduces memory usage by 40.47% and computation cost by 35.15% while providing comparable GLUE score. This is because SparseAdapter performs pruning of redundant parameters but uses unstructured pruning with masking, which does not actually improve resource efficiency.
@@
Figure 5 shows the freezing patterns of $\text{BERT}_{\text{large}}$ fine-tuned with SAFE — we observe similar patterns for other tasks. We find that SAFE tends to freeze adapters more in layers closer to the input layer. Such behavior aligns with our empirical observations presented in Figure 2 where adapters closer to the output layer need to be further adapted to the downstream tasks compared to those in earlier layers, contributing more to model performance.
@@ 5.2.2 Question Answering
5.2.3 Natural Language Generation
5.3 Regularization Effect
Table 3: Experimental results on natural language generation from the E2E NLG Challenge. For all metrics, higher is better. We report memory usage reduction in blue.
@@ 7 Limitations
We suggest the need for combination with prior research on memory-efficient training. These include low precision, microbatching, weight sharding, and gradient checkpointing techniques. Though we have not evaluated SAFE along with such memory-efficient training methods, SAFE can be complementarily used along with the methods since SAFE can be applied independently of the training method or weight precision. In particular, since the quantization-based compression technique is quite popular and effective in terms of both compression ratio and preservation of final accuracy, favorable results are expected from combining the proposed technique with the memory-efficient training methods Han et al. (2015).
@@ C.1 Model
We conduct experiments using a pre-trained model deployed on HuggingFace Wolf et al. (2019). For experiments on the NLU and QA benchmarks, we use bert-base-uncased and bert-large-uncased trained on BookCorpus, a dataset consisting of 11,038 unpublished books and English Wikipedia. We use roberta-base and roberta-large trained on 5 datasets (BookCorpus, English Wikipedia, CC-News, OpenWebText, and Stories) for the RoBERTa model. For experiments on the NLG benchmark, we use GPT-2 medium and GPT-2 large distributed by OpenAI. We use vit-base-patch16-224-in21k and vit-large-patch16-224-in21k distributed by Google for experiments in the ViT model. Finally, We use the swin-base-patch4-window7-224 and swin-large-patch4-window7-224 models distributed by Microsoft for experiments in the SWIN model.
@@ C.2 Computing Resources
Our experimental setup leverages 2 RTX4090 with 24GB memory for NLU, QA, and NLG tasks and 1 RTX 4090 for CV downstream task.
@@ C.3 Dataset Statistics
We present the dataset statistics of GLUE and SQuAD in following table.
Table 7: Summary of the NLP and QA benchmarks.
@@ C.4 Hyperparameter Settings
pre-trained model GPT-2 medium GPT-2 large
@@
final learning rate 1.00E-04 5.00E-05
batch size 8 4
# epochs 10 10
Seq Length 512 512

### v2

@@ Abstract
Transformer-based large-scale pre-trained models achieve great success. Fine-tuning is the standard practice for leveraging these models in downstream tasks. Among the fine-tuning methods, adapter-tuning provides a parameter-efficient fine-tuning by introducing lightweight trainable modules while keeping most pre-trained parameters frozen. However, existing adapter-tuning methods still impose substantial resource usage. Through our investigation, we show that each adapter unequally contributes to both task performance and resource usage. Motivated by this insight, we propose Selective Adapter FrEezing (SAFE), which gradually freezes less important adapters early to reduce unnecessary resource usage while maintaining performance. In our experiments, SAFE reduces memory usage, computation amount, and training time by 42.85%, 34.59%, and 11.82%, respectively, while achieving comparable or better task performance compared to the baseline. We also demonstrate that SAFE induces regularization effect, thereby smoothing the loss landscape, which enables the model to generalize better by avoiding sharp minima.
@@ 1 Introduction
Large-scale pre-trained language models (PLMs) have manifested superior performance in various tasks Kenton and Toutanova (2019); Liu et al. (2019); Radford et al. (2019); Yang et al. (2019). However, training PLMs from the scratch is time-consuming and resource-intensive. Common practice has been hence to fine-tune the large-scale pre-trained models by adapting all the parameters with the downstream tasks, i.e., full parameter fine-tuning (full-tuning).
Recently, Parameter-Efficient Fine-Tuning (PEFT), which focuses on optimizing a small fraction of parameters for downstream tasks, is receiving much attention Houlsby et al. (2019); Lester et al. (2021); Li and Liang (2021); Liu et al. (2022, 2023). Among various PEFT strategies, adapter-tuning has emerged as a prevalent method. It integrates lightweight modules, termed adapters, into each layer of PLMs and only tunes the adapters with the downstream tasks. As shown in Figure 1(a), the adapter-tuning methods Houlsby et al. (2019); Pfeiffer et al. (2021); Zaken et al. (2022); Hu et al. (2022); Zhang et al. (2022), significantly reduce the number of trainable parameters compared to the full-tuning while exhibiting better performance on a downstream task.
@@ 2 Related Work
Figure 2: (a) Accuracy and (b) Resource usage of adapter-tuning by injecting an adapter into each transformer layer of $\text{BERT}_{\text{base}}$ model on MNLI and QNLI datasets from GLUE.
@@
Resource Efficient Fine-Tuning: Several works have tried to target resource efficient fine-tuning. AdapterDrop Rücklé et al. (2021), randomly excludes partial adapters from each training step. However, it cannot de-allocate the activation memory for the adapters, because of the random selections — an adapter excluded from training in a step can be included in training in the following steps. SparseAdapterHe et al. (2022) applies unstructured pruning to the adapters. However, it also does not reduce the actual memory usage — this is because the weight matrices pruned with zero values still need to be fully allocated in the memory. LoRAPruneZhang et al. (2024) employs structured pruning for LoRA. Unfortunately, the aforementioned methods usually have an adverse impact on the accuracy. MEFT Liao et al. (2023) applies a reversible model to PEFT. By using the reversible network, MEFT calculates activations with accumulated outputs of layers, without saving the intermediate activations reducing the activation memory. However, calculations of the activations severely degrades the training time performance.
@@ 3 Motivation
Figure 3: Visualization of representation similarity between the trained model and the model at different training steps during adapter-tuning of the $\text{BERT}_{\text{base}}$ model on the MNLI and QNLI datasets from GLUE.
@@ 4.2 Adapter Freezing
In the freezing stage, SAFE gradually freezes adapters based on their importance score. At $t_{w}$-th epoch, SAFE compares the importance score of adapters with threshold $\tau_{T}$. If the importance score of an adapter is lower than $\tau_{T}$, SAFE identifies the adapter as a freezing candidate. After identifying freezing candidates, SAFE freezes them based on a moving threshold until $t_{f}$-th epoch — it increases the threshold from 0 to $\tau_{T}$ [^2] We empirically determine $\tau_{T}$ and final freezing epochs $t_{f}$ based on extensive experiments with various models and datasets. between $t_{w}$-th and $t_{f}$-th epochs following a cubic schedule as Zhang et al. (2022):
@@
where $t$ is the current epoch, $t_{w}$ is the number of initial warm-up epochs, $t_{f}$ is the number of final freezing epochs. The cubic schedule follows a non-linear growth pattern, enabling rapid exploration of the search space during initial stages and gradually slowing down to reliably converge to the target point. We leverage these advantages by freezing more less-important adapters in the early stages. As the number of trainable parameters in the model decreases, SAFE gradually reduces the number of freezing adapters, reaching the threshold $\tau_{T}$ stably.
@@ 4.3 Regularization Effect of SAFE
By selectively freezing less critical adapters, SAFE induces a regularization effect within the model. In transformer-based PLM $\mathcal{N}_{0}$, each of the $l$ transformer blocks $T_{l}$ is equipped with a distinct set of parameters $\theta^{0}_{l}$ for $l\in\{1,\ldots,n\}$. To reduce the computational overhead of fine-tuning all parameters $\theta^{0}_{l}$, lightweight adapters $\Delta\theta_{l}$ are introduced. To clarify how introducing adapters contribute to performance enhancements, Fu et al. (2023) formalize the optimization function as follows:
@@
where $\theta=\theta^{0}+M\Delta\theta$ and $M\in\{0,1\}^{m\times m}$, with $m=\text{dim}(\theta)$, serves as a diagonal matrix for selective parameter adjustment. Each diagonal element $M_{ii}\in\{0,1\}$ indicates whether the corresponding parameter of $\Delta\theta_{i}$ is active ($1$) or inactive (00), with all off-diagonal elements $M_{ij}$ set to 0. The regularization term is crucial for explaining how parameter constraints introduced by adapters can enhance model performance on downstream tasks. The $rank(M)$ is bounded by $m$, reflecting full capacity for parameter adaptation within each transformer block. However, such an approach can lead to excessive computation (See Figure 1). In contrast, our study explores the implications of constraining the $rank(M)$ to a reduced upper bound from $m$ to the number of trainable parameters following the proposed freezing algorithm in Section 4.2. For instance, in our motivational analysis in Section 3, where we limit the number of trainable parameters to one adapter per layer, the $rank(M)$ is bounded by $\frac{m}{l}$ by selectively activating $\Delta W_{l}$ for $T_{l}$. This constraint not only optimizes computational efficiency but also preserves the adaptability essential for superior performance on downstream tasks, as evidenced by our empirical results detailed in Section 5.3.
@@ 5.1 Experimental Setting
Baselines: To evaluate the effectiveness of SAFE, we benchmark against state-of-the-art PEFT method, LoRA Hu et al. (2022). We compare SAFE with five effective resource efficient fine-tuning methods, a previous work Zhang et al. , AdapterDrop Rücklé et al. (2021), SparseAdapter He et al. (2022), LoRAPrune Zhang et al. (2024), and MEFT Liao et al. (2023). SAFE’s performance is further compared with four other PEFT methods such as Houlsby Houlsby et al. (2019), Pfeiffer Pfeiffer et al. (2020), BitFit Zaken et al. (2022), and the adaptive method AdaLoRA Zhang et al. (2022) to demonstrate its versatility and applicability across different adapter-tuning frameworks (see Figure 1 and Appendix B for detailed results). Comprehensive details on the experimental setup and hyperparameters, such as training epochs and batch sizes, can be found in Appendix C.
@@ 5.2.1 Natural Language Understanding
Figure 5 shows the freezing patterns of $\text{BERT}_{\text{large}}$ fine-tuned with SAFE — we observe similar patterns for other tasks. We find that SAFE tends to freeze adapters more in layers closer to the input layer. Such behavior aligns with empirical observations presented in Figure 2 where adapters closer to the output layer need further adaptation to the downstream tasks compared to those in earlier layers, contributing more to model performance.
Compared to a previous work Zhang et al. , which attaches adapters only to the upper layers, SAFE achieves 2.24% higher GLUE score. This is because some adapters attached to lower layers keep contributing to the adaptation even in later training epochs, as shown in Figure 5 (e.g., 2nd and 5th adapters in QNLI task). Compared to AdapterDrop, SAFE provides up to 2.69% higher score (MRPC) while reducing memory usage by 49.47% (0.91% higher GLUE score and 40.47% reduced memory usage on average). This is because AdapterDrop reduces computation costs (FLOPs) by randomly dropping adapters for each step, whereas SAFE selectively freezes less-important adapters preserving critical adapters and thus achieving better performance. AdapterDrop also does not lower memory usage because the memory allocated to the dropped adapters cannot be de-allocated for the next step — adapters dropped in a step may not be dropped in the next step. Compared to SparseAdapter, SAFE reduces memory usage by 40.47% and computation cost by 35.15% while providing comparable GLUE score. This is because SparseAdapter performs pruning of redundant parameters but uses unstructured pruning with masking, which does not actually improve resource efficiency. On the other hand, although LoRAPrune is effective to reduce the memory usage, it severely degrades the accuracy as it fully eliminates less important adapter weights through structural pruning. As a result, SAFE achieves 3.68% higher GLUE score compared to LoRAPrune. MEFT reduces the memory usage at the cost of more than 2x of the FLOPs and training time. This is because MEFT applies a reversible model to LoRA, which reduces memory usage by not caching intermediate activations, but causes substantial computation overhead due to recomputation. Overall, SAFE strikes a better accuracy/memory efficiency/training-time performance trade-off compared to SOTA methods.
@@
[…]
@@ 5.2.2 Question Answering
Table 3: Experimental results on natural language generation from the E2E NLG Challenge. For all metrics, higher is better. Note that we report memory usage reduction in blue.
@@ 7 Limitations
We suggest the need for combination with prior research on memory-efficient training. These include low precision, micro-batching, weight sharding, and gradient checkpointing techniques. Though we have not evaluated SAFE along with such memory-efficient training methods, SAFE can be complementarily used along with the methods since SAFE can be applied independently of the training method or weight precision. In particular, since the quantization-based compression technique is quite popular and effective in terms of both compression ratio and preservation of final accuracy, favorable results are expected from combining the proposed technique with the memory-efficient training methods Han et al. (2015).
Acknowledgments
This work was supported in part by National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (RS-2023-00212711), Institute of Information & Communications Technology Planning & Evaluation(IITP)-ITRC(Information Technology Research Center) grant funded by the MSIT(IITP-2025-RS-2023-00260091), ICT Creative Consilience Program through IITP grant funded by the MSIT(IITP-2025-RS-2020-II201819), IITP grant funded by the MSIT (No. RS-2024-00398353, Development of Countermeasure Technologies for Generative AI Security Threats). We also thank the members of the Korea University Intelligent Computer Architecture & Systems research Lab for their useful comments and discussions, as well as the anonymous reviewers for their helpful feedback.
@@ C.1 Model
We conduct experiments using a pre-trained model deployed on HuggingFace Wolf et al. (2019). For experiments on the NLU and QA benchmarks, we use bert-base-uncased and bert-large-uncased trained on BookCorpus, a dataset consisting of 11,038 unpublished books and English Wikipedia. We use roberta-base and roberta-large trained on 5 datasets (BookCorpus, English Wikipedia, CC-News, OpenWebText, and Stories) for the RoBERTa model. For experiments on the NLG benchmark, we use GPT-2 medium and GPT-2 large distributed by OpenAI. We also employ the LLaMA-2 7B model, an open-weight large language model released by Meta, which has been trained on a mixture of publicly available and licensed datasets, including Common Crawl, C4, GitHub, Wikipedia, Project Gutenberg, ArXiv, and Stack Exchange. We use vit-base-patch16-224-in21k and vit-large-patch16-224-in21k distributed by Google for experiments in the ViT model. Finally, We use the swin-base-patch4-window7-224 and swin-large-patch4-window7-224 models distributed by Microsoft for experiments in the SWIN model.
@@ C.2 Computing Resources
Our experimental setup leverages 2 RTX4090 with 24GB memory for NLU, QA, and NLG tasks and 1 RTX 4090 for CV downstream task, excluding the LLaMA experiments.
@@ C.3 Dataset Statistics
[…]
Table 7: Summary of the NLU, QA, and NLG benchmarks.
@@ C.4 Hyperparameter Settings
Since the lower layers of the BERT model are more easily converted to fine-tuning than the upper layers, we extend the previous study Zhang et al. that froze most of the lower layers with adapter-tuning and experiment with a baseline that naively inserts adapters only into six layers. For LoRAPrune, we configure the pruning ratio to 0.5, pruning frequency to 10, and the pruning metric to ’lora,’ as these settings have shown effective results in prior work zhang2023loraprune. In the case of MEFT, for a fair comparison, we set the reversible layers to half of the total number of layers. We employ the $\text{MEFT}_{1}$ method, where the $\mathcal{F}$ architecture is the layer, from the three reversible transformation methods proposed in MEFT. For the hyperparameters of MEFT, such as $\lambda$ and $\beta$, we follow the values proposed in previous work Liao et al. (2023) to ensure consistency in the evaluation.
@@
pre-trained model GPT-2 medium GPT-2 large LLaMA-2 7B
@@
final learning rate 1.00E-04 5.00E-05 1.00E-04
batch size 8 4 128 (w/ micro batch size 2)
# epochs 10 10 3
Seq Length 512 512 512